OpenStudy (anonymous):

The diameter of a circle is 5 ft. What is the area of the circle? Use 3.14 for Pi.

OpenStudy (anonymous):

\[area = \pi r^2\] and \[r = diameter/ 2\] so first, find r by dividing the diameter by 2, then find the area by squaring r and multiplying by pi.

OpenStudy (anonymous):

\[A = \frac{ \pi d ^{2} }{ 4 }\] A=Area d=diameter

OpenStudy (anonymous):

19.6?

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Area of a circle is \[\pi r^2\].... And Radius is half of diameter... r=d/2 So, area= \[\pi d^2/4\]